OpenStudy (anonymous):

In which polygon is the sum of the measures of the interior angles 540°? A. pentagon B. nonagon C. heptagon D. hexagon

OpenStudy (midhun.madhu1987):

Use the formula... Total Interior angle = (n-2) * 180 540 = (n-2) * 180 Find the value of n which is the number of sides...

OpenStudy (anonymous):

got A thanks

OpenStudy (midhun.madhu1987):

yw :)